# Macdonald Index From Refined Kontsevich-Soibelman Operator

### v1

@@ Abstract
We propose a refinement of the Kontsevich-Soibelman operator for a class of “special” 4d $\mathcal{N}=2$ superconformal field theories characterized by the following conditions: (1) their Coulomb branch admits a source/sink chamber, i.e., a chamber in which the BPS quiver consists of only source and sink nodes, (2) The nodes with valency greater than 2 of the BPS quiver in a source/sink chamber are either all sources or all sinks. We present strong evidence that the trace of this refined operator is related to the Macdonald index of the theory. In particular, we conjecture closed form expressions for the Macdonald indices of the $(A_{1},\mathfrak{g})$ Argyres-Douglas theories for any simply-laced Lie algebra $\mathfrak{g}$.
@@ II.1 Trace Of KS Operator And Schur Index
Imposing the cyclicity of the trace, the trace of any element of $\cal A$} can then be written in terms of $q$ and $N_{f}$ variables $\text{Tr}[\gamma_{f_{i}}]$. These variables are functions of the flavor fugacities $z_{1},\dots,z_{n}$ and the explicit functional relation can be determined as explained in [Cordova:2015nma]. In [Cordova:2015nma], the trace of the KS operator was computed for a large class of examples and the conjecture (7) was verified.
@@ III.1 BPS Quiver And Source/Sink Chamber
The BPS quiver of a 4d ${\cal N}=2$ theory captures the spectrum of BPS states of the theory at a given point the Coulomb branch. The rules to draw the BPS quiver are as follows [Cecotti:2010fi, Alim:2011ae]:
@@ III.2 Refined KS Operator
$\begin{split}{\cal O}(q,T):=&\prod_{\gamma\in\text{source}}E_{q,T}((-1)^{n}q^{n/2}X_{-\gamma})^{(-1)^{n}\Omega_{n}(-\gamma)}\\ \times&\prod_{\gamma\in\text{sink}}\widetilde{E}_{q,T}((-1)^{n}q^{n/2}X_{-\gamma})^{(-1)^{n}\Omega_{n}(-\gamma)}\\ \times&\prod_{\gamma\in\text{source}}E_{q,T}((-1)^{n}q^{n/2}X_{\gamma})^{(-1)^{n}\Omega_{n}(\gamma)}\\ \times&\prod_{\gamma\in\text{sink}}E_{q}((-1)^{n}q^{n/2}X_{\gamma})^{(-1)^{n}\Omega_{n}(\gamma)}~.\end{split}$ (15)
@@
$\begin{split}{\cal O}(q,T):=&\prod_{\gamma\in\text{source}}\widetilde{E}_{q,T}((-1)^{n}q^{n/2}X_{-\gamma})^{(-1)^{n}\Omega_{n}(-\gamma)}\\ \times&\prod_{\gamma\in\text{sink}}E_{q,T}((-1)^{n}q^{n/2}X_{-\gamma})^{(-1)^{n}\Omega_{n}(-\gamma)}\\ \times&\prod_{\gamma\in\text{source}}E_{q}((-1)^{n}q^{n/2}X_{\gamma})^{(-1)^{n}\Omega_{n}(\gamma)}\\ \times&\prod_{\gamma\in\text{sink}}E_{q,T}((-1)^{n}q^{n/2}X_{\gamma})^{(-1)^{n}\Omega_{n}(\gamma)}~.\end{split}$ (16)
@@ IV.1.1 $(A_{1},A_{k})$ Theory
Figure 1: BPS quiver of the $(A_{1},A_{2n})$ theory in the source/sink chamber. Figure adapted from [Cordova:2015nma]
@@
Figure 2: BPS quiver of the $(A_{1},A_{2n+1})$ theory in the source/sink chamber. Figure adapted from [Cordova:2015nma]
@@ IV.1.2 $(A_{1},D_{k})$ Theory
Figure 3: BPS quiver of the $(A_{1},D_{2n+1})$ theory in the source/sink chamber. Figure adapted from [Cordova:2015nma]
Figure 4: BPS quiver of the $(A_{1},D_{2n+2})$ theory in the source/sink chamber. Figure adapted from [Cordova:2015nma]
@@
$\begin{split}&{{\cal I}}_{\text{M}}^{D_{2n+1}}(q,T,z)=(q)_{\infty}^{n}(Tq)_{\infty}^{n}\sum_{\begin{subarray}{c}\ell_{1},\cdots,\ell_{2n+1}\geq 0\\ k_{2n},k_{2n+1}\geq 0\end{subarray}}\frac{q^{\sum_{i=1}^{2n+1}\ell_{i}+\frac{1}{2}\sum_{i,j=1}^{2n+1}b^{D_{2n+1}}_{ij}\ell_{i}\ell_{j}}T^{\ell_{2n}+\ell_{2n+1}+\sum_{i=1}^{n-1}\ell_{2i-1}}z^{2(\ell_{2n+1}-k_{2n+1})}}{(q)_{k_{2n}}(q)_{k_{2n+1}}(q)_{\ell_{2n}}(q)_{\ell_{2n+1}}\prod_{i=1}^{n}(q)_{\ell_{2i-1}}^{2}\prod_{i=1}^{n-1}(q)_{\ell_{2i}}(Tq)_{\ell_{2i}}}\delta_{k_{2n}+k_{2n+1},\ell_{2n}+\ell_{2n+1}}~,\end{split}$ (25)
@@ IV.1.3 $(A_{1},E_{n})$ Theory
Figure 5: BPS quiver of the $(A_{1},E_{6})$ theory in the source/sink chamber. Figure adapted from [Cordova:2015nma]
@@
Figure 6: BPS quiver of the $(A_{1},E_{8})$ theory in the source/sink chamber. Figure adapted from [Cordova:2015nma]
@@
Figure 7: BPS quiver of the $(A_{1},E_{7})$ theory in the source/sink chamber. Figure adapted from [Cordova:2015nma]
@@ V Future Directions
Acknowledgments. R.K.S. would like to thank LITP, Stanford University for hospitality where this work began. The work of A.B. is supported by the US Department of Energy under grant DE-SC1019775. The work of R.K.S. is supported by the US Department of Energy under grant DE-SC0010008.

### v2

@@ Abstract
We propose a refinement of the Kontsevich-Soibelman operator for a class of “special” 4d $\mathcal{N}=2$ superconformal field theories characterized by the following conditions: (1) their Coulomb branch admits a source/sink chamber, i.e., a chamber in which the BPS quiver consists of only source and sink nodes, (2) The nodes with valency greater than 2 of the BPS quiver in a source/sink chamber are either all sources or all sinks. We present strong evidence that the trace of this refined operator is related to the Macdonald index of the theory. In particular, we conjecture new closed form expressions for the Macdonald indices of the $(A_{1},\mathfrak{g})$ Argyres-Douglas theories for any simply-laced Lie algebra $\mathfrak{g}$.
@@ I Introduction
Note Added. After we submitted our paper to the arxiv, the paper [Kim:2025klh] appeared which also conjectures expressions for the Macdonald indices of $(A_{1},\mathfrak{g})$ theories using the KS monodromy operator.
@@ II.1 Trace Of KS Operator And Schur Index
Imposing the cyclicity of the trace, the trace of any element of $\cal A$} can then be written in terms of $q$ and $N_{f}$ variables $\text{Tr}[X_{\gamma_{f_{i}}}]$. These variables are functions of the flavor fugacities $z_{1},\dots,z_{n}$ and the explicit functional relation can be determined as explained in [Cordova:2015nma]. In [Cordova:2015nma], the trace of the KS operator was computed for a large class of examples and the conjecture (7) was verified.
@@ III.1 BPS Quiver And Source/Sink Chamber
The BPS quiver of a 4d ${\cal N}=2$ theory [Denef:2002ru] captures the spectrum of BPS states of the theory at a given point the Coulomb branch. We will not go into relevance of quivers in 4d $\mathcal{N}=2$ theories in this work, and instead refer readers to [Denef:2002ru, Denef:2007vg, Manschot:2010qz, Manschot:2012rx] for more details. The rules to draw the BPS quiver are as follows [Cecotti:2010fi, Alim:2011ae]:
@@ III.2 Refined KS Operator
$\begin{split}{\cal O}(q,T):=&\prod_{\gamma\in\text{source}}\prod_{n\in\mathbb{Z}}E_{q,T}((-1)^{n}q^{n/2}X_{-\gamma})^{(-1)^{n}\Omega_{n}(-\gamma)}\\ \times&\prod_{\gamma\in\text{sink}}\prod_{n\in\mathbb{Z}}\widetilde{E}_{q,T}((-1)^{n}q^{n/2}X_{-\gamma})^{(-1)^{n}\Omega_{n}(-\gamma)}\\ \times&\prod_{\gamma\in\text{source}}\prod_{n\in\mathbb{Z}}E_{q,T}((-1)^{n}q^{n/2}X_{\gamma})^{(-1)^{n}\Omega_{n}(\gamma)}\\ \times&\prod_{\gamma\in\text{sink}}\prod_{n\in\mathbb{Z}}E_{q}((-1)^{n}q^{n/2}X_{\gamma})^{(-1)^{n}\Omega_{n}(\gamma)}~.\end{split}$ (15)
@@
$\begin{split}{\cal O}(q,T):=&\prod_{\gamma\in\text{source}}\prod_{n\in\mathbb{Z}}\widetilde{E}_{q,T}((-1)^{n}q^{n/2}X_{-\gamma})^{(-1)^{n}\Omega_{n}(-\gamma)}\\ \times&\prod_{\gamma\in\text{sink}}\prod_{n\in\mathbb{Z}}E_{q,T}((-1)^{n}q^{n/2}X_{-\gamma})^{(-1)^{n}\Omega_{n}(-\gamma)}\\ \times&\prod_{\gamma\in\text{source}}\prod_{n\in\mathbb{Z}}E_{q}((-1)^{n}q^{n/2}X_{\gamma})^{(-1)^{n}\Omega_{n}(\gamma)}\\ \times&\prod_{\gamma\in\text{sink}}\prod_{n\in\mathbb{Z}}E_{q,T}((-1)^{n}q^{n/2}X_{\gamma})^{(-1)^{n}\Omega_{n}(\gamma)}~.\end{split}$ (16)
@@ IV.1.1 $(A_{1},A_{k})$ Theory
Figure 1: BPS quiver of the $(A_{1},A_{2n})$ theory in the source/sink chamber. Figure adapted from [Cordova:2015nma].
@@
Figure 2: BPS quiver of the $(A_{1},A_{2n+1})$ theory in the source/sink chamber. Figure adapted from [Cordova:2015nma].
@@ IV.1.2 $(A_{1},D_{k})$ Theory
Figure 3: BPS quiver of the $(A_{1},D_{2n+1})$ theory in the source/sink chamber. Figure adapted from [Cordova:2015nma].
Figure 4: BPS quiver of the $(A_{1},D_{2n+2})$ theory in the source/sink chamber. Figure adapted from [Cordova:2015nma].
@@
$\begin{split}&{{\cal I}}_{\text{M}}^{D_{2n+1}}(q,T,z)=(q)_{\infty}^{n}(Tq)_{\infty}^{n}\sum_{\begin{subarray}{c}\ell_{1},\cdots,\ell_{2n+1}\geq 0\\ k_{2n},k_{2n+1}\geq 0\end{subarray}}\frac{q^{\sum_{i=1}^{2n+1}\ell_{i}+\frac{1}{2}\sum_{i,j=1}^{2n+1}b^{D_{2n+1}}_{ij}\ell_{i}\ell_{j}}T^{\ell_{2n}+\ell_{2n+1}+\sum_{i=1}^{n-1}\ell_{2i}}z^{2(\ell_{2n+1}-k_{2n+1})}}{(q)_{k_{2n}}(q)_{k_{2n+1}}(q)_{\ell_{2n}}(q)_{\ell_{2n+1}}\prod_{i=1}^{n-1}(q)_{\ell_{2i}}^{2}\prod_{i=1}^{n}(q)_{\ell_{2i-1}}(Tq)_{\ell_{2i-1}}}\delta_{k_{2n}+k_{2n+1},\ell_{2n}+\ell_{2n+1}}~,\end{split}$ (25)
@@ IV.1.3 $(A_{1},E_{n})$ Theory
Figure 5: BPS quiver of the $(A_{1},E_{6})$ theory in the source/sink chamber. Figure adapted from [Cordova:2015nma].
@@
Figure 6: BPS quiver of the $(A_{1},E_{8})$ theory in the source/sink chamber. Figure adapted from [Cordova:2015nma].
@@
Figure 7: BPS quiver of the $(A_{1},E_{7})$ theory in the source/sink chamber. Figure adapted from [Cordova:2015nma].
@@ V Future Directions
Acknowledgments. The authors would like to thank Gregory Moore, Leonardo Rastelli and Anirudh Deb for comments on the draft and Ashoke Sen for some useful discussions on related topics. R.K.S. would like to thank LITP, Stanford University for hospitality where this work began. The work of A.B. is supported by the US Department of Energy under grant DE-SC1019775. The work of R.K.S. is supported by the US Department of Energy under grant DE-SC0010008.